OpenStudy (anonymous):

Danielle's basic cell phone rate each month is $29.95. Add to that $5.95 for voice mail and $2.95 for text messaging. This past month Danielle spent an additional C dollars on long distance. Her total bill was $77.70. How much did Danielle spend on long distance?

OpenStudy (anonymous):

29.95+5.95+2.95+C=77.70 C=77.70-(29.95+5.95+2.95) C=77.70-38.85 C=38.85 She spent $38.85 on long distance, and she should probably switch to a flat rate plan. :)